OpenStudy (pphalke):

how much time is needed for $1,800 to accumulate $360 in simple interest at a rate of 10%?

OpenStudy (pphalke):

please help

OpenStudy (agent0smith):

What's the simple interest formula?

OpenStudy (pphalke):

SI=PRT

OpenStudy (agent0smith):

So, plug in your given numbers.

OpenStudy (pphalke):

its asking for how much time is needed, how do I put that in with other number

OpenStudy (agent0smith):

Plug in the numbers you have

OpenStudy (agent0smith):

Anything you don't have, leave as the letter

OpenStudy (pphalke):

how do I do that?

OpenStudy (pphalke):

anyone?

OpenStudy (chrishogan):

Okay first you need to find out what 10% of $1,800 is. Have and ideas on what it is?

OpenStudy (chrishogan):

any*

OpenStudy (agent0smith):

SI=PRT P is the principal amount, the 1800 R is the interest rate, but convert it to a decimal T is time, leave it as T SI is the interest, 360

OpenStudy (chrishogan):

$360=$180(T)

OpenStudy (pphalke):

@agent0smith it would be $360=180*0.1*t?

OpenStudy (agent0smith):

Close but P is 1800, not 180.

OpenStudy (pphalke):

hw do I solve this?

OpenStudy (pphalke):

@ChrisHogan

OpenStudy (chrishogan):

???

OpenStudy (agent0smith):

First, write out the equation correctly, since you made the mistake i pointed out.

OpenStudy (pphalke):

$360=1800*0.1*t?

OpenStudy (agent0smith):

Yes.

OpenStudy (pphalke):

how do I solve this to get the answer though?

OpenStudy (chrishogan):

did you get my message? @pphalke

OpenStudy (pphalke):

yes but @agent0smith had the equation set up right now how do I solve it

OpenStudy (agent0smith):

360=1800*0.1*t First simplify the right side. Then you can divide both sides

OpenStudy (pphalke):

how do I simplify that side

OpenStudy (agent0smith):

1800*0.1?

OpenStudy (pphalke):

its 180

OpenStudy (pphalke):

what do I do after I get this answer do I subtract something else

OpenStudy (agent0smith):

360=1800*0.1*t now we have 360=180*t divide both sides by 180